# Pre-training Large Memory Language Models with Internal and External Knowledge

### v2

@@ B.1 Evaluation Benchmarks
We evaluate factual precision using FactScore [36], a benchmark for open-ended biography generation. Given a generated text, FactScore extracts a set of atomic facts and computes the proportion that is supported by a trusted knowledge source. We use the first 100 biography queries provided in the benchmark. All models generate outputs using greedy decoding (maximum length = 256 tokens; repetition penalty = 1.2). Factuality is validated using retrieval-augmented prompting with ChatGPT, following the official evaluation protocol. [^4] https://github.com/shmsw25/FActScore
@@
We adapt the T-REx subset from LAMA [44] for autoregressive models by filtering examples in which the masked entity does not appear in the final position. This yields 11,615 left-to-right compatible examples, following Schick et al. [50]. Each input consists of a factual statement, such as: “Jaideep Sahni (born 1968) is an Indian [MASK]” The model is expected to complete the statement with a single token (e.g., actor). We evaluate using two metrics: Exact Match, which checks whether the reference answer appears among the first five generated content words, and Precision@1, which checks whether the first generated content token matches the reference.
@@ B.2 Machine Unlearning Setting
The TOFU benchmark [33] evaluates unlearning efficacy in privacy-sensitive scenarios, aiming to selectively remove a specific subset of information (the Forget Set) from a model (the Full Model) while maintaining performance on retained information (the Retain Set) and general model capabilities. The benchmark’s primary goal is to ensure the resulting Unlearned Model is statistically indistinguishable from a model trained exclusively on the Retain Set (the Retain Model).
@@
We use LLaMa-3.2-1B-Instruct as the base model and compare against NPO [60], a state-of-the-art method for unlearning. We test on forget 5% setting. For LmLm, unlearning is implemented simply by removing relevant entries corresponding to the Forget Set from the external database.
@@
Negative Preference Optimization (NPO) [60] is a state-of-the-art method for selective knowledge removal, especially effective for unlearning large portions (50%-90%) of training data. Unlike traditional gradient-based unlearning methods, which often degrade a model’s general performance and struggle to forget only 10% of training data, NPO explicitly guides the model away from undesired (negative) samples, maintaining stable training dynamics and preventing catastrophic performance degradation. Further details can be found in the original paper [60]. We follow the official TOFU implementation, running unlearning fine-tuning three times with different random seeds and reporting the mean and variance in Figure 6.
@@
For ROUGE evaluations, generated answers are post-processed to remove structured lookup tokens before computing the scores. For likelihood-based metrics (Answer Probability and Truth Ratio), we evaluate the model’s probabilities only on the annotated training input segments (_$\mathcal{I}\_\text{train}$, as defined in Appendix A.3). Thus, ROUGE scores remain directly comparable across methods, but likelihood-based metrics are not directly comparable due to differences between raw and annotated reference answers.
@@
B.3 Experimental Settings by Figure
@@
  • Figure 4: (Left) We eval on a held-out wikipedia validaition set of 100 passages (~21k tokens) every 1000 steps during pretraining. (Middle) The detailed results for FactScore and NLU are in Table 1 and  9. (Right) The backbone model is LLaMA-3.2-1B-Instruct. We implement LmLm evaluation based on TOFU official repo It is forget 5% setting. The detailed results for baselines methods are copied from TOFU repo.
@@ D.1 Does LmLm Affect Language Understanding?
Beyond the promising results, it is important to verify that our approach does not compromise the general capabilities of pretrained language models. To assess this, we follow Penedo et al. [41] and evaluate on a set of "high-signal" Natural Language Understanding (NLU) benchmarks (as shown in Table 9). Given our focus on smaller models, we exclude benchmarks where both GPT-2 and similarly sized models fail to rise above the noise floor [12]. This leaves us with the following benchmarks: Commonsense QA [54], HellaSwag [58], PIQA [4], SIQA [49], and ARC Easy [10]. Implementation details can be found in lighteval [^7] https://github.com/huggingface/lighteval.
@@ D.2 Entity Frequency Analysis in the Knowledge Database
We analyze the prevalence of knowledge triplets in our database by estimating how often each entity appears in Wikipedia. Using entity frequency statistics from Kandpal et al. [26], we apply fuzzy string matching (threshold = 70) to align database entities with Wikipedia entries. Figure 10 shows the distribution of matched entity frequencies. The results indicate that our database spans both common and long-tail knowledge, with a substantial portion of entities being less frequent in the overall training corpus.
@@ D.4 Qualitative Results
To complement our quantitative results, we present qualitative examples comparing outputs from LmLm, Standard, and off-the-shelf models. As shown in Table 11,  12 and 13, LmLm produces concise, factually grounded responses by leveraging external knowledge, whereas standard and off-the-shelf models often include hallucinated content. These examples illustrate how knowledge offloading enables LmLm to maintain factual precision, further supporting our central finding (Figure 4) that LmLms achieve more with less by scaling efficiently through externalized knowledge.
@@
LLaMA2-176M-Standard Ko Itakura is the name given to an ancient Japanese settlement in what is now part of Shimane Prefecture, Japan. The site was discovered by archaeologists from Kyoto University and has been excavated since 1978. The first written reference to this place appears in "Nihon Kōki", which dates back to the year 604 AD (Kumano no Kan’i), but it seems that there were already settlements at this location before then. In the early Heian period, the area around modern-day Shimonoseki City became known as , meaning "the land between rivers". During the Kamakura period, the region came under control of the Ashikaga clan who built their capital on Mount Hiei near present-day Naka Ward. After the fall of the Ashikaga shogunate, they lost control over much of eastern Japan, including the area surrounding modern-day Shimane prefecture. However, after the Battle of Sekigahara, Tokugawa Ieyasu established himself as lord of most of northern Japan with his domain centered on Matsue Castle. This castle later fell into disrepair during the Edo Period, and
@@
*The Artist’s Way* by Julia Cameron – Book Review | The Guardian
@@
Model Output (OpenAI/GPT2-355M) Kang Ji-hwan is known primarily for portraying characters with complex emotions like "the heartbreak," which can be seen throughout most drama shows nowadays. His character arc often involves both romantic relationships and difficult situations where people have conflicting feelings about their own lives. For example during this season we see how complicated it is for two men trying desperately not get divorced because there isn’t enough time left before graduation! What do you think? Do your favorite actors play these kinds/characters differently than others? Let us know what kind(ies?) I missed out below…
Model Output (Llama-3.2-1B-Instruct) Kang Ji-hwan is a renowned South Korean actor known for his versatility and range in various genres. Early Life Kang Ji-hwan was born on March 20, 1977, in Busan, South Korea. His full name is Jo Tae-gyu. Career Beginnings He started his career in musical theater before transitioning into film and television. Kang debuted in musical performances in 2001 with stints in productions like "The Rocky Horror Show" and "Grease". Breakthrough Roles His breakthrough came in 2005 with the drama "Be Strong, Geum-soon!". He later starred in "Capital Scandal", "Hong Gil-dong", and "90 Days, Time to Love", showcasing his ability to play complex characters. Film Career Kang ventured into cinema with movies like "Rough Cut" (2008) and "My Girlfriend Is an Agent" (2009). Recognition Throughout his career, Kang has received numerous awards and nominations, solidifying his status in Korean entertainment.

### v3

@@ B.1 Evaluation Benchmarks
We evaluate factual precision using FactScore (Min et al., 2023), a benchmark for open-ended biography generation. Given a generated text, FactScore extracts a set of atomic facts and computes the proportion that is supported by a trusted knowledge source. We use the first 100 biography queries provided in the benchmark. All models generate outputs using greedy decoding (maximum length = 256 tokens; repetition penalty = 1.2). Factuality is validated using retrieval-augmented prompting with ChatGPT, following the official evaluation protocol. [^5] https://github.com/shmsw25/FActScore
@@
We adapt the T-REx subset from LAMA (Petroni et al., 2019) for autoregressive models by filtering examples in which the masked entity does not appear in the final position. This yields 11,615 left-to-right compatible examples, following Schick et al. (2023). Each input consists of a factual statement, such as: “Jaideep Sahni (born 1968) is an Indian [MASK]” The model is expected to complete the statement with a single token (e.g., actor). We evaluate using two metrics: Exact Match, which checks whether the reference answer appears among the first five generated content words, and Precision@1, which checks whether the first generated content token matches the reference.
@@
PopQA.
We evaluate on the long-tail subset of PopQA, which contains 1,399 queries about rare entities (fewer than 100 monthly Wikipedia page views), following Asai et al. (2023). Performance is measured by Exact Match (EM), i.e., whether the gold answer appears in the model output. To enable fair comparison with small pre-trained models that are not instruction-tuned, we convert the QA format into a knowledge-completion task by prompting GPT-4 to rewrite each query. This reduces dependence on instruction-following ability and allows answers to be appended directly. All results reported in the paper use this rewritten format. For example: Original: What is Ufa the capital of? Rewritten: What is Ufa the capital of? Ufa is the capital of.
@@ B.2 Machine Unlearning Setting
The TOFU benchmark (Maini et al., 2024) evaluates unlearning efficacy in privacy-sensitive scenarios, aiming to selectively remove a specific subset of information (the Forget Set) from a model (the Full Model) while maintaining performance on retained information (the Retain Set) and general model capabilities. The benchmark’s primary goal is to ensure the resulting Unlearned Model is statistically indistinguishable from a model trained exclusively on the Retain Set (the Retain Model).
@@
We use LLaMa-3.2-1B-Instruct as the base model and compare against NPO (Zhang et al., 2024), a state-of-the-art method for unlearning. We test on forget 5% setting. For LmLm, unlearning is implemented simply by removing relevant entries corresponding to the Forget Set from the external database.
@@
Negative Preference Optimization (NPO) (Zhang et al., 2024) is a state-of-the-art method for selective knowledge removal, especially effective for unlearning large portions (50%-90%) of training data. Unlike traditional gradient-based unlearning methods, which often degrade a model’s general performance and struggle to forget only 10% of training data, NPO explicitly guides the model away from undesired (negative) samples, maintaining stable training dynamics and preventing catastrophic performance degradation. Further details can be found in the original paper (Zhang et al., 2024). We follow the official TOFU implementation, running unlearning fine-tuning three times with different random seeds and reporting the mean and variance in  Figure 6.
@@
For ROUGE evaluations, generated answers are post-processed to remove structured lookup tokens before computing the scores. For likelihood-based metrics (Answer Probability and Truth Ratio), we evaluate the model’s probabilities only on the annotated training input segments ($\mathcal{T}\_\text{train}$, as defined in Appendix A.3). Thus, ROUGE scores remain directly comparable across methods, but likelihood-based metrics are not directly comparable due to differences between raw and annotated reference answers.
@@
B.3 Experimental Setting of Further Analysis
Selective Knowledge Offloading.
We conduct a preliminary study on the trade-off between storing knowledge in model parameters and offloading it into the database. To enable selective externalization, we revert only a subset of annotations based on model learning difficulty.
Specifically, we first train a LmLm model with the original annotated data described in Sec. 4 and a Standard model with the unannotated data for one epoch. We measure the difference of the language modeling loss for the 5 tokens following each lookup for the LmLm and the corresponding tokens in the clean document for the Standard model. A larger gap suggests that Standard struggles to memorize the fact—typically long-tail or highly specific knowledge—making it a strong candidate for externalization. A smaller gap indicates that (i) the fact can be easily memorized by Standard, (ii) contextual hints make it trivial, or (iii) database lookups in LmLm are noisy or less useful. We calculate this loss difference across the full Wikipedia dataset, and include the distribution ( Figure 12) and random qualitative examples from each quantile bucket for illustration ( Table 14).
Based on this criterion, we then pretrain LLMs with varying knowledge offloading ratios $0\%,25\%,50\%,75\%,90\%,100\%]$}. Here, $0\%$ corresponds to Standard (fully parametric) and $100\%$ to LmLm (fully externalized). Offloading is applied progressively, starting with triplets that show the largest loss differences. The corresponding loss thresholds are: 3.51 for 10% offloading, 2.24 for 25%, 1.25 for 50%, 0.64 for 75%, and 0.13 for 90%.
B.4 Experimental Settings by Figure
@@
  • Figure 4: (Left) We eval on a held-out wikipedia validaition set of 100 passages (~21k tokens) every 1000 steps during pretraining. (Middle) The detailed results for FactScore and NLU are in  Table 2 and  12. (Right) The backbone model is LLaMA-3.2-1B-Instruct. We implement LmLm evaluation based on TOFU official repo It is forget 5% setting. The detailed results for baselines methods are copied from TOFU repo.
@@ Appendix C Detailed Results
Comparison with memory-based related work.
Table 10: Comparison of LmLm against prior memory-based work on EM (%). All baseline numbers are reported from their original papers. “Params” refers to non-embedding model parameters. “External Memory” summarizes the format or corpus scale used at inference.
Model Params External Memory T-REx EM (%)
CoLAKE 0.11B  5M Wikidata Triples 28.8
K-Adapter 0.34B  42M Adapter Parameters 29.1
BERT-KNN 0.34B  900M Instance Datastore 38.7
EaE 0.11B  1M Entity Memory 38.6
FILM 0.11B  1.7M Wikidata Triples 44.2
Toolformer 6.0B  API Access to Wikipedia 53.5
LmLm 0.16B  54.6M Triples 54.1
LmLm 0.36B  54.6M Triples 58.1
We also compare LmLm with prior memory-augmented language models ( Table 11). Several approaches externalize factual knowledge by attaching explicit memory modules. Fact-Injected Language Model (FILM) (Verga et al., 2021) augments LMs with a neuro-symbolic memory of entity–fact pairs. CoLAKE (Sun et al., 2020) jointly pretrains language and knowledge representations through a modified Transformer encoder. BERT-kNN (Kassner & Schütze, 2020) and EaE (Févry et al., 2020) use kNN-style retrieval or entity-specific memories to handle rare entities. K-Adapter (Wang et al., 2020) injects structured knowledge into PLMs via adapter modules. While these methods improve factuality by coupling models with external memories, they generally do not address controllable unlearning, which distinguishes LmLm from this line of work.
Table 10 shows that LmLm achieves consistently higher precision (Exact Match) on T-REx compared to prior memory-augmented and tool-augmented models, while using fewer parameters and without depending on large external tool APIs. These results highlight both the modeling advantages of our approach and its empirical gains over existing memory-based baselines.
Table 11: Comprehensive comparison of related methods. = supported, = not supported, and $\sim$ = partial support (e.g., edit without unlearning, or not explicitly evaluated in prior work).
Model Architecture Knowledge Storage Performance Goal
[…]
Pretraining
REALM (Guu et al., 2020) Enc-only Docs Prompt prepend Joint retriever + LM
RETRO (Borgeaud et al., 2022) Dec-only Docs Cross-attn. to retrieved chunks Scale with retrieval
Atlas (Izacard et al., 2022) Enc–dec Docs FiD (passages) Few-shot generalization
Memorizing Transformer (Wu et al., 2022) Dec-only (params+cache) kNN attention (cache) Long-term memory
MemSinks (Ghosal et al., 2025) Dec-only (sink neurons) Memorization isolation Localize memorization
LMLM (ours) Dec-only (partial) KB Explicit lookup Decouple facts from weights
Post-training
K-Adapter (Wang et al., 2020) Enc + adapters (params+adapt.) Adapter layers $\sim$ Inject KB via adapters
RAG (Lewis et al., 2021) Enc–dec Docs Retrieved docs in prompt Improve factuality
Toolformer (Schick et al., 2023) Dec-only APIs Tool calls Extend via tools
MEMIT (Meng et al., 2023b) Dec-only Weight editing Direct fact edit
Inference
kNN-LM (Shi et al., 2022) Dec-only DS kNN search + prob. interp. Neighbor interpolation
RAG-variant Dec-only Docs Retrieved docs in prompt LLM + retrieval
Memory-based Models
CoLAKE (Sun et al., 2020) Enc-only KG Joint MLM over KG Joint language + knowledge repr.
BERT-kNN (Kassner & Schütze, 2020) Enc-only Docs kNN search QA for long-tail facts via retrieval
EaE (Févry et al., 2020) Enc-only Entity Mem Entity Mem + prob. interp. Entity-specific memories
FILM (Verga et al., 2021) Enc-only KB Entity-fact Mem + prob. interp. $\sim$ Neuro-symbolic KB with fact injection
@@ D.1 Does LmLm Affect Language Understanding?
Beyond the promising results, it is important to verify that our approach does not compromise the general capabilities of pretrained language models. To assess this, we follow Penedo et al. (2024) and evaluate on a set of ”high-signal” Natural Language Understanding (NLU) benchmarks (as shown in  Table 12). Given our focus on smaller models, we exclude benchmarks where both GPT-2 and similarly sized models fail to rise above the noise floor (Du et al., 2025). This leaves us with the following benchmarks: Commonsense QA (Talmor et al., 2019), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2019), SIQA (Sap et al., 2019), and ARC Easy (Clark et al., 2018). Implementation details can be found in lighteval [^8] https://github.com/huggingface/lighteval.
@@ D.2 Entity Frequency Analysis in the Knowledge Database
We analyze the prevalence of knowledge triplets in our database by estimating how often each entity appears in Wikipedia. Using entity frequency statistics from Kandpal et al. (2023), we apply fuzzy string matching (threshold = 70) to align database entities with Wikipedia entries.  Figure 11 shows the distribution of matched entity frequencies. The results indicate that our database spans both common and long-tail knowledge, with a substantial portion of entities being less frequent in the overall training corpus.
@@
Figure 12: Distribution of per-triplet loss differences with quantile buckets used to select triplets for externalization.
@@ D.4 Qualitative Results
To complement our quantitative results, we present qualitative examples comparing outputs from LmLm, Standard, and off-the-shelf models. As shown in  Table 15,  16 and 17, LmLm produces concise, factually grounded responses by leveraging external knowledge, whereas standard and off-the-shelf models often include hallucinated content. These examples illustrate how knowledge offloading enables LmLm to maintain factual precision, further supporting our central finding ( Figure 4) that LmLms achieve more with less by scaling efficiently through externalized knowledge.
Table 14: Qualitative examples of triplets selected for externalization at different quantile buckets of per-triplet loss difference. Higher quantiles correspond to facts that are harder for Standard to memorize (long-tail or specific knowledge), while lower quantiles correspond to easier or contextually obvious facts. The triplet of interest in each example is highlighted in color.
Delta Loss Quantile Example
3.31 75%-100% Joseph Charles “Joe” Avellone III, M.D. (born <|db_start|> Joseph Charles Avellone III<|sep|> Birth Date<|db_retrieve|> September 29, 1948<|db_end|> September 29, 1948) is an <|db_start|> Joseph Charles Avellone III<|sep|> Nationality<|db_retrieve|> American<|db_end|> American <|db_start|> Joseph Charles Avellone III<|sep|> Occupation<|db_retrieve|> medical doctor, businessman, politician<|db_end|> medical doctor, businessman, and politician from <|db_start|> Joseph Charles Avellone III<|sep|> State of Residence<|db_retrieve|> Massachusetts<|db_end|> Massachusetts…He then worked as CEO of biomedical company <|db_start|> Joseph Charles Avellone III<|sep|> CEO Of<|db_retrieve|> Veritas Medicine<|db_end|> Veritas Medicine
3.23 75%-100% Elsa van Dien…Her thesis, supervised by <|db_start|> Elsa van Dien<|sep|> Thesis Supervisor<|db_retrieve|> Donald Menzel<|db_end|> Donald Menzel, discussed the Stark effect in the Balmer lines of early type stars.
1.41 50%-75% Kermia albicaudata is a <|db_start|> Kermia albicaudata<|sep|> Species Type<|db_retrieve|> sea snail<|db_end|> species of sea snail, a marine gastropod mollusk in the <|db_start|> Kermia albicaudata<|sep|> Family<|db_retrieve|> Raphitomidae<|db_end|> family Raphitomidae…. Distribution. This species occurs in the <|db_start|> Kermia albicaudata<|sep|> Geographical Distribution<|db_retrieve|> Persian Gulf<|db_end|> Persian Gulf.
0.82 25%-50% Havering Country Park… The land was purchased by the <|db_start|> Havering Country Park<|sep|> Purchased By<|db_retrieve|> Greater London Council<|db_end|> Greater London Council and opened to the public in <|db_start|> Havering Country Park<|sep|> Opening Year<|db_retrieve|> 1975<|db_end|> 1975, with ownership transferring to <|db_start|> Havering Country Park<|sep|> Ownership Transferred To<|db_retrieve|> Havering Council<|db_end|> Havering Council in <|db_start|> Havering Council<|sep|> Ownership Year<|db_retrieve|> 1986<|db_end|> 1986.
0.003 0%-25% Kagoshima District, Kagoshima is a district located in <|db_start|> Kagoshima District, Kagoshima<|sep|> Located In Prefecture<|db_retrieve|> Kagoshima Prefecture, Japan<|db_end|> Kagoshima Prefecture, Japan.
@@
LLaMA2-176M-Standard Ko Itakura is the name given to an ancient Japanese settlement in what is now part of Shimane Prefecture, Japan. The site was discovered by archaeologists from Kyoto University and has been excavated since 1978. The first written reference to this place appears in ”Nihon Kōki”, which dates back to the year 604 AD (Kumano no Kan’i), but it seems that there were already settlements at this location before then. In the early Heian period, the area around modern-day Shimonoseki City became known as , meaning ”the land between rivers”. During the Kamakura period, the region came under control of the Ashikaga clan who built their capital on Mount Hiei near present-day Naka Ward. After the fall of the Ashikaga shogunate, they lost control over much of eastern Japan, including the area surrounding modern-day Shimane prefecture. However, after the Battle of Sekigahara, Tokugawa Ieyasu established himself as lord of most of northern Japan with his domain centered on Matsue Castle. This castle later fell into disrepair during the Edo Period, and
@@
*The Artist’s Way* by Julia Cameron – Book Review — The Guardian
@@
Model Output (OpenAI/GPT2-355M) Kang Ji-hwan is known primarily for portraying characters with complex emotions like ”the heartbreak,” which can be seen throughout most drama shows nowadays. His character arc often involves both romantic relationships and difficult situations where people have conflicting feelings about their own lives. For example during this season we see how complicated it is for two men trying desperately not get divorced because there isn’t enough time left before graduation! What do you think? Do your favorite actors play these kinds/characters differently than others? Let us know what kind(ies?) I missed out below…
Model Output (Llama-3.2-1B-Instruct) Kang Ji-hwan is a renowned South Korean actor known for his versatility and range in various genres. Early Life Kang Ji-hwan was born on March 20, 1977, in Busan, South Korea. His full name is Jo Tae-gyu. Career Beginnings He started his career in musical theater before transitioning into film and television. Kang debuted in musical performances in 2001 with stints in productions like ”The Rocky Horror Show” and ”Grease”. Breakthrough Roles His breakthrough came in 2005 with the drama ”Be Strong, Geum-soon!”. He later starred in ”Capital Scandal”, ”Hong Gil-dong”, and ”90 Days, Time to Love”, showcasing his ability to play complex characters. Film Career Kang ventured into cinema with movies like ”Rough Cut” (2008) and ”My Girlfriend Is an Agent” (2009). Recognition Throughout his career, Kang has received numerous awards and nominations, solidifying his status in Korean entertainment.
Appendix E LLM Usage
We used LLMs as general-purpose assistive tools to improve clarity, grammar, and readability of the paper. LLMs were also used for creating tables and figures, but not for research ideation, experiment design, or substantive writing. All research contributions, methodology, experiments, and analyses were conceived, implemented, and written by the authors.